### v1

@@ 4.2 Example 2: Slightly extended source
Based on the above information, we created a mock dataset which exhibits energy-dependent morphology. This consisted of simulating a 3D dataset with a livetime of 50 hours based on the Cherenkov Telescope Array Observatory (CTAO) IRFs (Cherenkov Telescope Array Observatory & Cherenkov Telescope Array Consortium, 2021), assuming a specific source model. The source is slightly extended as described by a circular emission region which follows a power law distribution $\sigma=\sigma_{0}E^{-\alpha}$, where $\sigma_{0}=0.1^{\circ}$, $\alpha=0.3$ and $E$ is the energy axis for the given geometry. The source has a flux of a few percent of the Crab flux at 1 TeV ($4.22\times 10^{-13}$ cm-2 s-1 TeV-1) and an index of 2.42, mimicking the gamma-ray source HESS J1833--105 (a composite SNR and PWN system, H. E. S. S. Collaboration et al., 2018; Djannati-Ataï et al., 2008). These models are assigned to the dataset to predict the counts data based on a Poisson probability distribution. To test for the energy-dependent morphology using the tool, we chose three energy bins for this example: 0.2-2 TeV, 2-10 TeV and 10-100 TeV.
@@
Figure 4: The results of the estimator tool for the mock dataset. The different coloured rings show the fitted GaussianSpatialModel in the various energy bins along with their uncertainties. The plus symbols show the position and error from the fit.
@@
Figure 5: Comparison of the variation of the extension with energy. The results obtained from the estimator for example 2 are shown in blue and the expected extension values taken from the simulated source sizes in grey. The grey line is described by a power law of $\sigma_{0}E^{-\alpha}$, where $\sigma_{0}=0.1^{\circ}$ and $\alpha=0.3$.
@@ 5 Validations
As previously mentioned, Wilks’ theorem (Wilks, 1938) states that the test statistic ($\Delta$TS) follows a chi-squared distribution for a large number of samples. To validate the accuracy of our method, we need to conduct tests to ensure this holds true. We utilise a bootstrap parametric method in which the same method as the energy-dependent tool is used. We generate a simulated Poisson distributed dataset based on the model of interest for the chosen source. This mock dataset consists of 10 hours of observations based on the CTAO IRFs (Cherenkov Telescope Array Observatory & Cherenkov Telescope Array Consortium, 2021) and exhibits clear energy-dependent morphology. This simulation is performed 1000 times, resulting in a large sample set that contains the various parameters of interest including $\Delta$TS, as well as the latitude, longitude and sigma for each sample. Figure 6 shows how the test statistic is distributed compared to the expected chi-squared function, with six degrees of freedom corresponding to the difference in the number of free parameters between the null and alternative hypotheses, as described in section 2.
Figure 6: Distribution of the samples of $\Delta$TS, calculated through Equation 1, for each of the 1000 simulations. The expected chi-squared function has six degrees of freedom, which is the difference between the number of free parameters in the global fit (H0) versus the individual fits (H1).
@@
Figure 7 shows the distribution of the weighted chi-squared values for each of the free spatial parameters (longitude, latitude and sigma) calculated via Equation 3 for each sample set. The results follow a $\chi^{2}$ distribution, with the number of degrees of freedom corresponding to the number of energy bands minus one.
Figure 7: Distribution of $\chi_{p}^{2}$ values as calculated through Equation 3 for each of the free parameters in the energy-dependent estimator tool.
The distributions in Figure 6 and Figure 7 show good agreement with Equation 3, especially where the sample density is highest. The deviations between the simulated distribution and the corresponding line equation, especially in the tail regions, are likely attributed to having an insufficient number of samples. Insufficient sampling can also lead to statistical fluctuations, resulting in decreased precision therefore making it harder for the simulated data to perfectly match the theoretical expectation. To test the compatibility of these we utilised a Kolmogorov-Smirnov test for goodness of fit. Typically a threshold p-value of 0.05 is utilised to indicate the data is compatible with the expected function. Comparison between the $\Delta$TS and expected chi-squared function in Figure 6 leads to a p-value=0.097, supporting the notion that there is no significant difference between the two. We perform the same test for Figure 7 on the $\chi_{p}^{2}$ values, assuming they follow the expected chi-squared function with two degrees of freedom, and find a p-value$>0.05$ for each, providing evidence of compatibility of the two.
The $\chi_{p}^{2}$ values for each parameter (as shown in Figure 7) are calculated using Equation 3. For each sample set, the three $\chi_{p}^{2}$ values (for sigma, longitude and latitude) are compared and the largest value is taken as the ‘$\max(\chi_{p}^{2})$’. It is expected that these maximum values follow a specific distribution function:
$P=N\,f(x,df)(F(x,df))^{N-1}$ (4)
where $N$ is the number of parameters, $f(x,df)$ is the probability density function (PDF) and $F(x,df)$ is the cumulative density function (CDF). Here, $x$ are the $\max(\chi_{p}^{2})$ values and $df$ are the number of energy bands minus one. The derivation of Equation 4 is shown in Appendix B. The distribution of the $\max(\chi_{p}^{2})$ values, along with the expected function (Equation 4) are shown in Figure 8.
Figure 8: Distribution of the $\max(\chi_{p}^{2})$ value (either longitude, latitude or sigma) for each simulation.
The next step involves the comparison of the two distinct methods for calculating the significance, namely $\Delta$TS and $\chi^{2}$. This allows us to determine the compatibility of the results obtained from the two methods. As both the $\Delta$TS and max$(\chi_{p}^{2})$ functions follow a $\chi^{2}$ distribution, we can directly compare their relationship on a scatter plot, as presented in the top panel of Figure 9. It is worth reiterating that the $\Delta$TS method is the recommended approach, as it accounts for correlations between parameters, in contrast to the $\chi_{p}^{2}$ statistic, which treats each parameter independently. As a result, we do not necessarily expect a one-to-one relationship between the two. We can also compare the p-values derived from each method. The equation for the p-value corresponding to the $\max(\chi_{p}^{2})$ values is defined as follows:
@@
where $F$ is the cumulative density function which gives the probability of the random variable $X$ being less than or equal to $x$. The number of degrees of freedom, $df$, is equal to the number of energy bands minus one and $N$ is the number of parameters of interest, in this case equal to three.
The p-value for the $\Delta$TS is calculated through:
$p_{\rm{value},\Delta\rm{TS}}=s(\Delta\mathrm{TS},df)$ (6)
where $s$ is the survival function (also defined as $1-F$, where $F$ denotes the CDF). The number of degrees of freedom from the $\Delta$TS calculations is given by $df$, which in this case is equal to six. The comparison of the p-value from these methods are shown in the middle panel of Figure 9.
Finally, we can also determine the expected significance from each of the aforementioned methods through the following equations. For the $\max(\chi_{p}^{2})$ values the significance is calculated through:
$\sigma_{\chi^{2}}=s^{-1}(0.5*p_{\rm{value},\chi_{p}^{2}})$ (7)
The significance from the $\Delta$TS value is given by:
@@
where $s^{-1}$ is the inverse survival function (inverse of $s$). The comparison of the distribution of significance for each method is shown in the bottom panel of Figure 9. The middle and bottom panels of Figure 9 include a black dashed line corresponding to unity ($x=y$), serving as a visual reference for assessing agreement between the two distributions.
Figure 9: Comparison of the $\Delta$TS and $\chi^{2}$ methods. The black dashed line in the middle and bottom panel show the line of unity ($x=y$). Top: direct comparison of $\Delta$TS (Equation 1) with the $\max(\chi_{p}^{2})$ value (Equation 3) from the three parameters (longitude, latitude or sigma) for each simulation. Middle: the p-values for the $\Delta$TS and $\max(\chi_{p}^{2})$ method as calculated through Equation 6 and Equation 5, respectively. Bottom: the significance for the $\Delta$TS and $\max(\chi_{p}^{2})$ method as calculated through Equation 8 and Equation 7, respectively.
@@ 6 Conclusion
We assess the accuracy of our technique by conducting a number of tests on a mock dataset that exhibits energy-dependent morphology. We utilise 1000 simulations to ensure the method is robust. In general, we find that these results are in good agreement with those derived from established statistical equations and methods.
@@ Appendix A Code Example Output
Figure 10: Output from the code example shown in Figure 1.
Appendix B Chi-squared Maximums Derivation
For a single chi-squared measurement:
@@ Appendix B Chi-squared Maximums Derivation
where $f(x,df)$ is the chi-squared probability density function (PDF) and $df$ is the number of degrees of freedom. For any $x$ to be lower than $X$, the probability is:
$P(\forall x_{i},x_{i}<X)=\prod_{i}\int_{X}^{\infty}f(x,df)$ (10)
Therefore, the probability that at least one $x$ is larger than $X$ is:
$\begin{split}P(\exists i,x_{i}>X)&=1-P(\forall x_{i},x_{i}<X)\\ &=1-\left(\int_{X}^{\infty}f(x,df)\right)^{N}\end{split}$ (11)
This is explicitly the cumulative density function (CDF). To obtain the PDF, we have to differentiate the CDF (Equation 11, leading to:
$\begin{split}P(x,df)&=Nf(x,df)\left(\int_{X}^{\infty}f(x,df)\right)^{(N-1)}\\ &=Nf(x,df)\left(F(x,df)\right)^{(N-1)}\end{split}$ (12)

### v3

@@ 4.2 Example 2: Slightly extended source
Based on the above information, we created a mock dataset that exhibits energy-dependent morphology. This consisted of simulating a 3D dataset with a live time of 50 hours based on the Cherenkov Telescope Array Observatory (CTAO) IRFs (Cherenkov Telescope Array Observatory & Cherenkov Telescope Array Consortium, 2021), assuming a specific source model. The source is slightly extended as described by a circular emission region that follows a power law distribution $\sigma=\sigma_{0}E^{-\alpha}$, where $\sigma_{0}=0.1^{\circ}$, $\alpha=0.3$ and $E$ is the energy axis for the given geometry. The source has a flux of a few percent of the Crab flux at 1 TeV ($4.22\times 10^{-13}$ cm-2 s-1 TeV-1) and an index of 2.42, mimicking the gamma-ray source HESS J1833--105 (a composite SNR and PWN system, H. E. S. S. Collaboration et al., 2018; Djannati-Ata¨ı et al., 2008). These models are assigned to the dataset to predict the counts data based on a Poisson probability distribution. To test for the energy-dependent morphology using the tool, we chose three energy bins for this example: 0.2-2 TeV, 2-10 TeV and 10-100 TeV.
@@
Figure 4: Results of the estimator tool for the mock dataset. The different coloured rings show the fitted GaussianSpatialModel in the various energy bins along with their uncertainties. The plus symbols show the position and error from the fit.
@@
Figure 5: Comparison of the variation in the extension with energy. The results obtained from the estimator for example 2 are shown in blue and the expected extension values taken from the simulated source sizes in grey. The grey line is described by a power law of $\sigma_{0}E^{-\alpha}$, where $\sigma_{0}=0.1^{\circ}$ and $\alpha=0.3$.
@@ 5 Validations
As previously mentioned, Wilks’ theorem (Wilks, 1938) states that the $\Delta$TS follows a chi-squared distribution for a large number of samples. To validate the accuracy of our method, we needed to conduct tests to ensure this holds true. We utilised a bootstrap parametric method in which the same method as the energy-dependent tool was used. We generated a simulated Poisson distributed dataset based on the model of interest for the chosen source. This mock dataset consisted of 10 hours of observations based on the CTAO IRFs (Cherenkov Telescope Array Observatory & Cherenkov Telescope Array Consortium, 2021) and exhibits a clear energy-dependent morphology. This simulation was performed 1000 times, resulting in a large sample set that contains the various parameters of interest including $\Delta$TS, as well as the latitude, longitude, and sigma for each sample. Figure 6 shows how the $\Delta\rm{TS}$ is distributed compared to the expected chi-squared function, with six degrees of freedom corresponding to the difference in the number of free parameters between the null and alternative hypotheses, as described in section 2.
Figure 6: Distribution of the samples of $\Delta$TS, calculated through Eq. 1, for each of the 1000 simulations. The expected chi-squared function has six degrees of freedom, which is the difference between the number of free parameters in the global fit (H0) versus the individual fits (H1).
@@
Figure 7 shows the distribution of the weighted chi-squared values for each of the free spatial parameters (longitude, latitude and sigma) calculated via Eq. 3 for each sample set. The results follow a $\chi^{2}$ distribution, with the number of degrees of freedom corresponding to the number of energy bands minus one.
Figure 7: Distribution of $\chi_{p}^{2}$ values as calculated through Eq. 3 for each of the free parameters in the energy-dependent estimator tool.
The distributions in Figs. 6 and 7 show good agreement with Eq. 3, especially where the sample density is highest. The deviations between the simulated distribution and the corresponding line equation, especially in the tail regions, are likely attributed to having an insufficient number of samples. Insufficient sampling can also lead to statistical fluctuations, resulting in decreased precision therefore making it harder for the simulated data to perfectly match the theoretical expectation. To test the compatibility of these we utilised a Kolmogorov-Smirnov test for goodness of fit. Typically a threshold p-value of 0.05 is utilised to indicate the data are compatible with the expected function. Comparison between the $\Delta$TS and expected chi-squared function in Fig. 6 leads to a p-value=0.097, supporting the notion that there is no significant difference between the two. We perform the same test for Fig. 7 on the $\chi_{p}^{2}$ values, assuming they follow the expected chi-squared function with two degrees of freedom, and find a p-value>0.05>0.05> 0.05 for each, providing evidence that the two are compatible.
The $\chi_{p}^{2}$ values for each parameter (as shown in Fig. 7) are calculated using Eq. 3. For each sample set, the three $\chi_{p}^{2}$ values (for sigma, longitude and latitude) are compared and the largest value is taken as the ‘$\max(\chi_{p}^{2})$’. It is expected that these maximum values follow a specific distribution function:
$P=N\,f(x,df)(F(x,df))^{N-1},$ (4)
where $N$ is the number of parameters, $f(x,df)$ is the probability density function (PDF) and $F(x,df)$ is the cumulative density function (CDF). Here, $x$ are the $\max(\chi_{p}^{2})$ values and $df$ are the number of energy bands minus one. The derivation of Eq. 4 is shown in Sect. B. The distribution of the $\max(\chi_{p}^{2})$ values, along with the expected function (Eq. 4) are shown in Fig. 8.
Figure 8: Distribution of the $\max(\chi_{p}^{2})$ value (longitude, latitude, or sigma) for each simulation.
[…]
@@
where $F$ is the CDF that gives the probability of the random variable $X$ being less than or equal to $x$. The number of degrees of freedom, $df$, is equal to the number of energy bands minus one and $N$ is the number of parameters of interest, in this case equal to three.
The p-value for the $\Delta$TS is calculated through
$p_{\rm{value},\Delta\rm{TS}}=s(\Delta\mathrm{TS},df),$ (6)
[…]
Finally, we can also determine the expected significance from each of the aforementioned methods through the following equations. For the $\max(\chi_{p}^{2})$ values the significance is calculated through
$\sigma_{\chi^{2}}=s^{-1}(0.5*p_{\rm{value},\chi_{p}^{2}}).$ (7)
The significance from the $\Delta$TS value is given by
@@
where $s^{-1}$ is the inverse survival function (inverse of $s$). The comparison of the distribution of significance for each method is shown in the bottom panel of Fig. 9. The middle and bottom panels of Fig.9 include a black dashed line corresponding to unity ($x=y$), serving as a visual reference for assessing agreement between the two distributions.
Figure 9: Comparison of the $\Delta$TS and $\chi^{2}$ methods. Top: Direct comparison of $\Delta$TS (Eq. 1) with the $\max(\chi_{p}^{2})$ value (Eq. 3) from the three parameters (longitude, latitude, or sigma) for each simulation. Middle: p-values for the $\Delta$TS and $\max(\chi_{p}^{2})$ method as calculated through Eqs. 6 and 5, respectively. Bottom: Significance for the $\Delta$TS and $\max(\chi_{p}^{2})$ method as calculated through Eqs. 8 and 7, respectively. The dashed black line in the middle and bottom panels shows the line of unity ($x=y$).
@@ 6 Conclusion
We assessed the accuracy of our technique by conducting a number of tests on a mock dataset that exhibits an energy-dependent morphology. We utilised 1000 simulations to ensure the method is robust. In general, we find that these results are in good agreement with those derived from established statistical equations and methods.
@@ Appendix A Code example output
Figure 10: Output from the code example shown in Fig. 1.
Appendix B Chi-squared maximum derivations
For a single chi-squared measurement,
@@ Appendix B Chi-squared maximum derivations
where $f(x,df)$ is the chi-squared PDF and $df$ is the number of degrees of freedom. For any $x$ to be lower than $X$, the probability is
$P(\forall x_{i},x_{i}<X)=\prod_{i}\int_{X}^{\infty}f(x,df).$ (10)
Therefore, the probability that at least one $x$ is larger than $X$ is
$\begin{split}P(\exists i,x_{i}>X)&=1-P(\forall x_{i},x_{i}<X)\\ &=1-\left(\int_{X}^{\infty}f(x,df)\right)^{N}\end{split}.$ (11)
This is explicitly the CDF. To obtain the PDF, we had to differentiate the CDF (Eq. 11), which led to
$\begin{split}P(x,df)&=Nf(x,df)\left(\int_{X}^{\infty}f(x,df)\right)^{(N-1)}\\ &=Nf(x,df)\left(F(x,df)\right)^{(N-1)}\end{split}.$ (12)